OpenStudy (anonymous):

A corner store sells two kinds of baked goods: cakes and pies. A cake costs $5 and a pie costs $7. In one day, the store sold 15 baked goods for a total of $91. How many cakes did they sell?

OpenStudy (kropot72):

Let the quantity of cakes sold = C Let the quantity of pies sold = P 5C + 7P = 91 C + P = 15 Multiply both sides of C + P = 15 by 7 to give: 7C + 7P = 105 Subtract 5C + 7P = 91 from the last equation to get: 2C = 14 C = 7 = quantity of cakes sold